# Mining Legal Arguments to Study Judicial Formalism

### v1

@@ Abstract
Courts must justify their decisions, but systematically analyzing judicial reasoning at scale remains difficult. This study refutes claims about formalistic judging in Central and Eastern Europe (CEE) by developing automated methods to detect and classify judicial reasoning in Czech Supreme Courts’ decisions using state-of-the-art natural language processing methods. We create the MADON dataset of 272 decisions from two Czech Supreme Courts with expert annotations of 9,183 paragraphs with eight argument types and holistic formalism labels for supervised training and evaluation. Using a corpus of 300k Czech court decisions, we adapt transformer LLMs for Czech legal domain by continued pretraining and experiment with methods to address dataset imbalance including asymmetric loss and class weighting. The best models successfully detect argumentative paragraphs (82.6% macro-F1), classify traditional types of legal argument (77.5% macro-F1), and classify decisions as formalistic/non-formalistic (83.2% macro-F1). Our three-stage pipeline combining ModernBERT, Llama 3.1, and traditional feature-based machine learning achieves promising results for decision classification while reducing computational costs and increasing explainability. Empirically, we challenge prevailing narratives about CEE formalism. This work shows that legal argument mining enables reliable judicial philosophy classification and shows the potential of legal argument mining for other important tasks in computational legal studies. Our methodology is easily replicable across jurisdictions, and our entire pipeline, datasets, guidelines, models, and source codes are available at https://github.com/trusthlt/madon/.
@@ 1 Introduction
Courts are expected to justify their decisions. How courts reason matters to scholars, practitioners, and citizens alike [Walton_Macagno_Sartor_2021]. Yet systematically analyzing judicial reasoning at scale remains difficult. It is a challenging interdisciplinary task that inherits the key problems from legal NLP, legal theory, and empirical legal studies. Empirical work often depends on manual annotations, which are time-consuming and costly [hall2008, gilardi_chatgpt_2023, braun_i_2023-1]. Legal concepts like “formalism” or traditional argument types remain disputed or abstract, hard to measure across large numbers of cases and foreign to NLP community [luders_classifying_2024-1, habernal_mining_2024-1, schauer_1988]. While recent advances in NLP and legal argument mining promise to support quantitative computational analysis of legal reasoning, key challenges persist: clarifying complex legal concepts [habernal_mining_2024-1, luders_classifying_2024-1], addressing scarce, imbalanced annotations with different quality [braun_i_2023-1], handling non-English legal texts, and ensuring explainable, reliable pipelines [Chau_Livermore_2024]. In short, judicial reasoning matters, but studying it at scale remains hard.
@@
We turn to Central and Eastern Europe (CEE), a region rarely examined empirically, yet widely alleged to be formalistic. Scholarship often portrays CEE judiciaries as rooted in late-communist legal culture and narrowly focused on text and procedure [cserne_discourses_2020-1, kuhn_judiciary_2011-2, matczak_eu_2015, matczak_constitutions_2010-2, manko_weeds_2013, bystranowski_formalist_2022-1, Bencze_2021], but much of this is anecdotal. We still know little about how CEE courts in fact justify their decisions [cserne_discourses_2020-1, cserne_judicial_2024, komarek_2015], and, to our knowledge, no study has used LLMs to test whether formalism indeed prevails in the region. Czechia offers an ideal case study: the Supreme Court (SC), largely inherited from the communist era, and the Supreme Administrative Court (SAC), founded in 2003 and seen as more reformist, are routinely described as opposites—rigidly formalistic SC versus more discursive SAC [matczak_eu_2015, kosar_domestic_2020-1, sipulova_kosar_purging_2024-1, kuhn_worlds_2004, kuhn_application_2005, kuhn_judiciary_2011-2, kuhn_quality_2018, matczak_constitutions_2010-2, Kadlec2016, stehlik_unijni_2014]. We use this setting to test the broader claims about allegedly omnipresent formalism in CEE as well as to evaluate the potential of legal NLP on non-English judgments.
@@
We ask three research questions, namely 1) what are the capabilities of state-of-the-art NLP methods to detect and classify traditional types of legal arguments in non-english judicial decisions, 2) can they distinguish between formalistic and non-formalistic reasoning styles at the decision level, and 3) does empirical analysis support or refute prevailing claims about CEE courts?
@@
By addressing these research questions, we make five substantial contributions to the field. First, we create the MADON dataset: a collection of 272 Czech Supreme Court and Supreme Administrative Court decisions with ca 9,183 paragraphs anno- tated by legal experts with eight argument types and one holistic formalism label. Second, we publish detailed annotation guidelines grounded in established theories of legal argumentation and refined through over 200 pilot cases. Third, we compile and preprocess a corpus of 300,000 anonymized Czech court decisions for continued pretraining of LLMs. Fourth, we experiment with pretraining, fine-tuning and testing different configurations of state-of-the-art open models (Llama 3.1, ModernBERT) across three tasks: detecting arguments, classifying argument types, and predicting formalism. From these results, we propose a unified pipeline to classify decisions as non/formalistic that integrates all tasks and model families, improving explainability, reducing computational costs, and boosting performance. Finally, we use the MADON dataset for first empirical analysis of the two Czech apex courts. In sum, we show how contemporary NLP research helps to investigate judicial philosophies and traditional jurisprudential questions; we provide novel data, pipelines, datasets and models for such investigations.
@@ 2.1 Jurisprudential debate on formalism
A simple example illustrates formalistic judging. Suppose a statute fixes a 15-day deadline to file an appeal. If a party files on day 16 because the counsel was hospitalized, a formalistic judge would dismiss the appeal as inadmissible, since the statute contains no exception. A non-formalistic judge would accept the appeal, invoking access-to-justice principles, the purpose of the rule, and the unfairness of harsh consequences under the individual circumstances.
Formalism concerns how lawyers, typically judges, approach legal questions. It is an approach to law and judicial adjudication that prioritizes the binding authority of legal norms and their wording. A formalist sees law as distinct from politics and limits the influence of non-legal norms and extra-legal considerations in answering legal questions, such as social, economical, political, or moral factors [weber_1922, schauer_1988, Weinrib1988].
@@
Formalism can be approached analytically, descriptively, or normatively. The analytical approach asks, e.g., what formalism is and whether it is possible (e.g. can a judge simply apply a rule given the very same legal system already contains open-textured principles). The descriptive approach asks whether judges in fact decide cases in a formalistic manner (e.g. would an average judge indeed reject the appeal on strictly procedural grounds?). The normative approach asks whether formalism is desirable and what judges should do when deciding cases - is it right that judges stick to statutory wording rather than purpose?
@@ 2.2 Legal formalism in Central and Eastern Europe (CEE)
Both descriptive and normative approaches toward formalism are especially prominent in the CEE.
Descriptively, formalism is widely portrayed as the dominant judicial style in the CEE region [bystranowski_formalist_2022-1, dixon_2023, Bencze_2021]. Existing scholarship traces the formalism to socialist-era legal culture and unreformed post-communist judiciary [cserne_discourses_2020-1, kosar_comparative_2023, manko_weeds_2013, Suteu_2023]. Labels include “hyperpositivism” [manko_weeds_2013], “mechanical jurisprudence” [kosar_comparative_2023], and “vulgar formalism” [Suteu_2023]. This portrayal appears across the region, in Poland [Kustra-Rogatka_2023, manko_weeds_2013], Czechia [kuhn_judiciary_2011-2, kosar_domestic_2020-1, Jakab_Dyevre_Itzcovich_2017], Romania [Suteu_2023], Serbia [Besirevic_2014], and parts of the post-Yugoslav space [uzelac_survival_2010]. It leads some theorists to endorse very strong claims suggesting that CEE legal systems might collectively form a distinct ”third legal family” alongside common and civil law traditions [manko_2016]. All this because the CEE judiciary is allegedly formalistic.
@@
From the normative perspective, formalism is heavily criticized in the CEE region. Scholars argue that it impedes rule of law [matczak_2018] and the enforcement of EU law [matczak_constitutions_2010-2, kuhn_worlds_2004]. According to kuhn_judiciary_2011-2, the formalism shows that the transition to liberal democracy was weak or incomplete, given the alleged communist-era roots of formalism. These criticisms are echoed in practice. The Czech Constitutional Court has repeatedly criticized the rest of the judiciary for excessive formalism and struck excessive formalism down as a violation of fundamental rights [constitutionalcourt2012]. Alleged formalism has also been raised in judicial appointments to constitutional courts as a reason to nominate more liberal judges [CzechSenate2002]. Beyond the scholarship and courts, politicians have critiqued formalism too: the Slovak President recently blamed judicial formalism for public frustration in post-communist Slovakia and Czechia [CT24_2024_Caputova]. Formalism has become central to debates about the region’s communist past and democratic future. Key actors view formalism as a toxic post-communist legacy undermining judicial credibility and the rule of law in the new democratic state.
@@
The formalism debate resonates particularly strongly in Czechia. There, it often concerns the country’s two supreme courts: the Supreme Court (SC), which handles civil and criminal matters, and the Supreme Administrative Court (SAC), deciding administrative law matters. The SC is characterized as an unreformed formalistic institution tied to the communist past, similar to most other courts in the judicial system [kuhn_worlds_2004, kuhn_application_2005, kuhn_judiciary_2011-2, kuhn_quality_2018, matczak_constitutions_2010-2, matczak_eu_2015, kosar_domestic_2020-1, sipulova_kosar_purging_2024-1, Kadlec2016, stehlik_unijni_2014]. Indeed, the SC was “inherited” from the communist era and staffed with many older judges who were formerly members of the communist party [sipulova_kosar_purging_2024-1]. By contrast, the newly established Supreme Administrative Court (SAC), created in 2003 to adjudicate administrative matters and reportedly staffed by new generation of jurists, is described as “discursive” and more open in its reasoning [matczak_eu_2015]. For over two decades, scholars have raised the distinct genealogies of these two courts as evidence of their divergent decision-making styles [kuhn_worlds_2004, kuhn_application_2005, kuhn_judiciary_2011-2, kuhn_quality_2018, matczak_constitutions_2010-2, matczak_eu_2015, kosar_domestic_2020-1, sipulova_kosar_purging_2024-1, Kadlec2016, stehlik_unijni_2014]. We call this the Tale of Two Courts.
Despite the omnipresent debates about formalism, however, little is known about how CEE judiciaries actually decide cases. One trend in the CEE literature [komarek_2015, cserne_judicial_2024, manko_weeds_2013] rightly emphasizes that claims of pervasive formalism are mostly anecdotal and based on the views of a few elite insiders, rather than systematic evidence [cserne_discourses_2020-1]. Without reliable data about judicial reasoning in the region, prevailing claims about CEE courts remain unsubstantiated. This applies equally to the Tale of Two Courts itself as there is no empirical study on the adjudication of SC and its reasoning. As [cserne_judicial_2024] concludes, empirical analysis of formalism in CEE requires clear definitions of formalism and remains in its early stages. Simply put, the evidence is missing.
@@
This paper aims to close the existing research gap by a pioneering empirical research in this direction.
2.3 Empirical analysis of formalism
We now discuss existing studies that investigate formalism empirically—both within the CEE region and beyond.
@@ 2.3 Empirical analysis of formalism
Three main approaches can be distinguished. The first approach studies the types of arguments courts employ. It quantifies the frequency of various argument types (e.g., linguistic, systematic, teleological) that are then classified as either formalistic (e.g., linguistic arguments) or non-formalistic (e.g., teleological arguments). This approach is not new; legal theory has investigated argument types for centuries, and typically US scholars like choi_2020 or Krishnakumar2020, krishnakumar_textualism_2024 have been empirically investigating argument types to assess prevailing interpretative theories. Fewer studies focus on argument types to measure formalism specifically. Most of these concern CEE judiciaries [matczak_constitutions_2010-2, matczak_eu_2015, bystranowski_formalist_2022-1]. The underlying assumption of this approach is that the more formalistic arguments a court uses, the more formalistic the court is.
@@
The second approach [stiglitz_h_historical_2024, thalken_modeling_2023] directly labels paragraphs as formalistic or non-formalistic without analyzing specific argument types. This simplification enabled Stiglitz and Thalken not only to manually code decisions but also to provide large-scale analysis using NLP. This approach assumes that the higher the number of formalistic paragraphs appearing in courts’ decisions, the more formalistic the court is.
@@
The third approach captures formalism’s different dimensions by examining various parameters simultaneously, including the court’s rhetorical style and procedural choices alongside argumentation patterns. This approach offers a holistic perspective on how formalism manifests across different dimensions of judicial decision-making [alberstein_sarat_measuring_2012, alberstein_whats_2020].
Despite valuable contributions, current approaches to measuring legal formalism face several problems. Most rely on small, manually coded samples without NLP, and none use argument mining to measure formalism directly. They typically cover only narrow slices of case law—e.g., a handful of criminal cases [bystranowski_formalist_2022-1] or selected administrative cases [matczak_constitutions_2010-2, matczak_eu_2015]. More importantly, many suffer from methodological shortcomings: missing or insufficient coding manuals and a lack of reported inter-annotator agreement, which undermines reproducibility [matczak_eu_2015, bystranowski_formalist_2022-1].
Each of the three approaches has additional limitations. First, the argument type approach can be insufficient: simple counts often miss crucial nuances. While this method yields useful quantitative data about which argument types courts use and how often, it can mislead as a measure of formalism when, for example, non-formalistic arguments are typically decisive but scarce while formalistic ones appear very often but are merely supplementary or peripheral. Simply put, a single well-put policy argument may outweigh multiple brief references to the wording, when assessing formalistic nature of a decision.
Second, the paragraph labeling approach likewise downplays the importance of particular passages. It also downplays argument types, the standard lens in legal theory and legal argumentation [Walton_Macagno_Sartor_2021, alexy_argument_2010] and provides less informative data on reasoning practices.
Finally, the multiparameter approach, though comprehensive, brings significant complexity to annotation.
@@
We differentiate substantially from previous works and combine the best of all worlds. First, we analyze argument types and go further by publishing detailed annotation guidelines, reporting inter-annotator agreement, and employing state of the art LLMs. Second, we combine fine-grained argument mining classifying paragraphs into 8 argument types with a binary decision level label formalistic/non-formalistic; to this end, we design unique pipeline that reduces computational requirements, increases explainability and boosts performance. Third, we consider multiple aspects of formalism as the last approach, but condense them into a single binary holistic label, formalistic or non formalistic, and also employ NLP to predict this variable. As such, our approach builds on existing methods, but enables more comprehensive measurement of judicial reasoning and formalism.
The discussed shortcomings of existing approaches show the need for methods that remain true to legal theory but can work on a larger scale. Legal argument mining offers such a path.

### v2

@@ Abstract
Courts must justify their decisions, but systematically analyzing judicial reasoning at scale remains difficult. This study tests claims about formalistic judging in Central and Eastern Europe (CEE) by developing automated methods to detect and classify judicial reasoning in decisions of Czech Supreme Courts using state-of-the-art natural language processing methods. We create the MADON dataset of 272 decisions from two Czech Supreme Courts with expert annotations of 9,183 paragraphs with eight argument types and holistic formalism labels for supervised training and evaluation. Using a corpus of 300,511 Czech court decisions, we adapt transformer LLMs to Czech legal domain through continued pretraining and we experiment with methods to address dataset imbalance including asymmetric loss and class weighting. The best models can detect argumentative paragraphs (82.6% Bal-F1), classify traditional types of legal argument (77.5% Bal-F1), and classify decisions as formalistic/non-formalistic (83.8% Bal-F1). Our three-stage pipeline combining ModernBERT, Llama 3.1, and traditional feature-based machine learning achieves promising results for decision classification while reducing computational costs and increasing explainability. Empirically, we challenge prevailing narratives about CEE formalism. We demonstrate that legal argument mining enables promising judicial philosophy classification and highlight its potential for other important tasks in computational legal studies. Our methodology can be used across jurisdictions, and our entire pipeline, datasets, guidelines, models, and source codes are available at https://github.com/trusthlt/madon.
@@ 1 Introduction
Courts are expected to justify their decisions. How courts reason matters to scholars, practitioners, and citizens alike [Walton_Macagno_Sartor_2021]. Yet systematically analyzing judicial reasoning at scale remains difficult. It is a challenging interdisciplinary task that inherits the key problems from legal NLP, legal theory, and empirical legal studies. Empirical work often depends on manual annotations, which are time-consuming and costly [hall2008, gilardi_chatgpt_2023, braun_i_2023-1]. Legal concepts like “formalism” or traditional argument types remain disputed or abstract, hard to measure across large numbers of cases and foreign to the NLP community [luders_classifying_2024-1, Habernal.et.al.2023.AILaw, schauer_1988]. While recent advances in NLP and legal argument mining promise to support quantitative computational analysis of legal reasoning, key challenges persist: clarifying complex legal concepts [Habernal.et.al.2023.AILaw, luders_classifying_2024-1], addressing scarce, imbalanced annotations with different quality [braun_i_2023-1], handling non-English legal texts, and ensuring explainable, reliable pipelines [Chau_Livermore_2024]. In short, judicial reasoning matters, but studying it at scale remains hard.
@@
We turn to Central and Eastern Europe (CEE), a region rarely examined empirically, yet widely alleged to be formalistic. Scholarship often portrays CEE judiciaries as rooted in late-communist legal culture and narrowly focused on text and procedure [cserne_discourses_2020-1, kuhn_judiciary_2011-2, matczak_eu_2015, matczak_constitutions_2010-2, manko_weeds_2013, bystranowski_formalist_2022-1, Bencze_2021]. Scholars describe courts in the region as exhibiting “hyperpositivism” [manko_weeds_2013], “mechanical jurisprudence” [Kosar2023], and “vulgar” or “deeply rooted formalism” [Suteu_2023, sipulova_kosar_purging_2024-1]. Formalism allegedly appears across the region, including Poland [Kustra-Rogatka_2023, manko_weeds_2013], Czechia [kuhn_judiciary_2011-2, kosar_domestic_2020-1, Jakab_Dyevre_Itzcovich_2017], Hungary [matczak_constitutions_2010-2], Romania [Suteu_2023], and Serbia [Besirevic_2014]. But much of this is anecdotal. We still know little about how CEE courts in fact justify their decisions [cserne_discourses_2020-1, cserne_judicial_2024].
@@
We ask three research questions. First, can state-of-the-art NLP methods detect argumentative paragraphs and classify them into eight traditional legal argument types in Czech judicial decisions? Second, can models distinguish between formalistic and non-formalistic reasoning styles at the decision level? And third, does an empirical analysis of legal reasoning in the Czech Supreme Court and Supreme Administrative Court support or challenge prevailing claims about formalism in Czech apex courts?
@@
By addressing these research questions, we make five substantial contributions to the field. First, we create the MADON dataset, a collection of 272 Czech Supreme Court and Supreme Administrative Court decisions with 9,183 paragraphs annotated by legal experts with eight argument types and one holistic formalism label. Second, we publish detailed annotation guidelines grounded in established theories of legal argumentation and refined through over 200 pilot cases. Third, we compile and preprocess a corpus of 300,511 anonymized Czech court decisions for continued pretraining of LLMs. Fourth, we experiment with pretraining, fine-tuning and testing different configurations of state-of-the-art and traditional open models (Llama 3.1, ModernBERT, MLP) across three tasks: detecting arguments, classifying argument types, and predicting formalism. From these results, we propose a unified pipeline to classify decisions as formalistic/non-formalistic that integrates all tasks and model families, improving explainability, reducing computational costs, and providing the best performance. Finally, we use the MADON dataset for the first empirical analysis of the two Czech apex courts. In sum, we show how contemporary NLP research helps to investigate judicial philosophies and traditional jurisprudential questions; we provide novel datasets, pipelines, and models for such investigations.
@@ 2.1 Jurisprudential debate on formalism
The following simple example illustrates formalistic judging. Suppose a statute fixes a 15-day deadline to file an appeal. If a party files on day 16 because the counsel was hospitalized, a formalistic judge would dismiss the appeal as inadmissible, since the statute contains no exception. A non-formalistic judge would accept the appeal, invoking access-to-justice principles, the purpose of the rule, and the unfairness of harsh consequences under the individual circumstances.
Formalism concerns how lawyers, typically judges, might approach legal questions. It is an approach to law and judicial adjudication that prioritizes the binding authority of legal norms and their wording. A formalist sees law as distinct from politics and limits the influence of non-legal norms and extra-legal considerations in answering legal questions, such as social, economic, political, or moral factors [weber_1922, schauer_1988, Weinrib1988].
@@
Formalism can be approached analytically, normatively, or descriptively. The analytical approach asks, e.g., what formalism is and whether it is possible (e.g., can a rule restrict a judge given the very same legal system already contains open-textured principles). The normative approach asks whether formalism is a proper way to deal with legal questions—is it right that judges stick to statutory wording rather than purpose? Lastly, the descriptive approach investigates whether judges in fact decide cases in a formalistic manner (e.g., would an average judge indeed stick to the letter of the law and disregard general values?). For example, schauer_1988 combines analytical and normative approach to clarify what formalism is and why it is desirable. An illustrative example of a descriptive approach would be that of stiglitz_h_historical_2024 who convincingly investigate how the formalism of US Supreme Court changed throughout history or whether formalistic judges stay formalist [Stiglitz2026]. The debate on CEE combines the descriptive and normative aspects by claiming that formalism is omnipresent and undesirable. We introduce this debate in the next section.
@@ 2.2 Legal formalism in Central and Eastern Europe (CEE)
Descriptively, formalism is widely portrayed as the dominant judicial style in the CEE region [bystranowski_formalist_2022-1, dixon_2023, Bencze_2021, kuhn_judiciary_2011-2, Kustra-Rogatka_2023, sipulova_kosar_purging_2024-1]. Some authors even suggest that the socialist past and formalist present might justify classifying Central European legal systems into a distinct “Central European Legal Family” [MankoCercelSulikowski2016, Manko2013_Legal_tradition]. In short, existing literature sees CEE judiciaries as distinct and formalistic.
@@
From the normative perspective, formalism is heavily criticized in the CEE region. Scholars argue that it impedes rule of law [matczak_2018] and the enforcement of EU law [matczak_constitutions_2010-2, kuhn_worlds_2004]. According to kuhn_judiciary_2011-2, the formalism shows that the transition to liberal democracy was weak or incomplete, given the alleged communist-era roots of formalism. These criticisms are echoed in practice. The Czech Constitutional Court has repeatedly criticized the rest of the judiciary for excessive formalism and struck excessive formalism down as a violation of fundamental rights [constitutionalcourt2010]. [^1] The court held: “The Constitutional Court has repeatedly demonstrated in its consistent case law that it does not tolerate, on the part of public authorities, and in particular the ordinary courts, a formalistic approach that employs sophisticated reasoning to enforce an evident injustice. (…) Excessive formalism in the interpretation of legal norms leading to an extremely unjust conclusion constitutes a violation of fundamental rights.” Alleged formalism has also been raised in judicial appointments to constitutional courts as a reason to nominate more liberal judges [CzechSenate2002]. Beyond the scholarship and courts, politicians have critiqued formalism too: the Slovak President recently blamed judicial formalism for public frustration in post-communist Slovakia and Czechia [CT24_2024_Caputova]. Formalism has become central to debates about the region’s communist past and democratic future. Key actors view formalism as a toxic post-communist legacy undermining judicial credibility and the rule of law in the new democratic state.
@@
The formalism debate resonates particularly strongly in Czechia, where it concerns the country’s two supreme courts: the Supreme Court (SC), handling civil and criminal matters, and the Supreme Administrative Court (SAC), deciding administrative matters. The SC is characterized as an unreformed formalistic institution inherited from the communist era [kuhn_worlds_2004, kuhn_application_2005, kuhn_judiciary_2011-2, kuhn_quality_2018, matczak_constitutions_2010-2, matczak_eu_2015, kosar_domestic_2020-1, sipulova_kosar_purging_2024-1, Kadlec2016, stehlik_unijni_2014]. By contrast, the SAC, created in 2003 and staffed by a new generation of jurists, is described as “discursive” and more open in its reasoning [matczak_eu_2015]. Allegedly, the SC is a formalistic post-communist court, while the SAC is not.
Despite the omnipresent debates, little is known about how CEE judiciaries actually decide cases, including the SC and SAC. Some of the recent scholarship rightly notes that the formalism narrative lacks clear definitions and remains mostly anecdotal [cserne_judicial_2024, cserne_discourses_2020-1, komarek_2015, manko_weeds_2013]. The evidence is missing.
@@
This paper aims to close the existing research gap by a pioneering empirical research in this direction. We next summarize current approaches to empirical analysis of formalism and how we build on them.
2.3 Existing approaches to empirical analysis of formalism
The literature offers three approaches to the empirical analysis of formalism.
@@ 2.3 Existing approaches to empirical analysis of formalism
The first approach studies the types of arguments courts employ. It quantifies the frequency of various argument types (e.g., linguistic, systemic, teleological) that are then classified as either formalistic (e.g., linguistic arguments) or non-formalistic (e.g., teleological arguments). This approach is not new; legal theory has investigated argument types for centuries, and typically US scholars like choi_2020 or Krishnakumar2020, krishnakumar_textualism_2024 have been empirically investigating argument types to assess prevailing interpretative theories. Fewer studies focus on argument types to measure formalism specifically. Most of these concern CEE judiciaries [matczak_constitutions_2010-2, matczak_eu_2015, bystranowski_formalist_2022-1]. The underlying assumption of this approach is that the more formalistic arguments a court uses, the more formalistic the court is.
@@
The second approach [Stiglitz2026, stiglitz_h_historical_2024, thalken_modeling_2023] directly labels paragraphs as formalistic or non-formalistic without analyzing specific argument types. This simplification enabled them not only to manually code decisions but also to provide large-scale analysis of US Supreme Court using NLP. This approach assumes that the higher the number of formalistic paragraphs appearing in courts’ decisions, the more formalistic the court is.
@@
The third approach captures different dimensions of formalism by examining various parameters simultaneously, including the court’s rhetorical style and procedural choices alongside argumentation patterns. This approach offers a holistic perspective on how formalism manifests across different dimensions of judicial decision-making and has been applied mainly in an Israeli context [alberstein_sarat_measuring_2012, alberstein_whats_2020].
Current approaches face several limitations. Most rely on small, manually coded samples, cover only narrow slices of case law, and have reproducibility issues as they do not publish annotation manuals or report inter-annotator agreement [matczak_eu_2015, bystranowski_formalist_2022-1]. None use argument mining to measure formalism directly. Each approach also has specific weaknesses: simply counting formalistic argument or paragraphs can mislead when a single decisive principle outweighs numerous peripheral textual references; paragraph-level labeling downplays argument types, the standard lens in legal theory [Walton_Macagno_Sartor_2021, alexy_argument_2010] and provides less informative data on reasoning practices in detail; and the multiparameter approach, though comprehensive, brings significant complexity to annotation.
@@
We combine the best of all three approaches. We analyze argument types with detailed annotation guidelines and reported inter-annotator agreement; we combine fine-grained argument mining across eight types with a binary holistic formalism label with a NLP pipeline that reduces computational requirements and increases explainability; and we condense multiple dimensions of formalism into a single binary label while employing NLP to predict it. Our approach thus enables richer measurement of judicial reasoning by making use of recent advances in legal argument mining, which we introduce next.